1. If a die is rolled 3 times, find the probability of getting at least 1 even number.

2. A student takes a 3-question, multiple-choice exam with 4 choices for each question out of which one choice is correct and guesses on each question. Find the probability that the student guessed two questions correctly.
Answer #1

1) P(even number) = 3/6 = 1/2

P(getting at least one even number) = 1 - P(no even number on three roll) = 1 - (1/2)3 = 7/8 = 0.875

2) P(guessing 1 question correctly) = 1/4 = 0.25

P(guessed 2 questions correctly) = 3C2 * 0.252 * (1 - 0.25)1 = 0.1406